# Seeing What’s Not There: Spurious Correlation in Multimodal LLMs

### v1

@@ 4 SpurLens: Automating Discovery of Spurious Cues
$\textbf{PA}_{s}$ $=\frac{1}{K}\sum\nolimits_{\mathcal{I}\in\mathcal{U}_{t,f_{i}}^{+}}\textbf{PA}(\mathcal{M},\mathcal{I},t)$ (11)
$\textbf{PA}_{c}$ $=\frac{1}{K}\sum\nolimits_{\mathcal{I}\in\mathcal{U}_{t,f_{i}}^{-}}\textbf{PA}(\mathcal{M},\mathcal{I},t)$ (12)
PA Gap $=\textbf{PA}_{s}-\textbf{PA}_{c}$ (13)
In the object recognition case, the Spuriosity Gap measures the difference in recognition accuracy between images with and without $f_{i}$, based on the top-$K$ and bottom-$K$ ranked images in the $f_{i}$-score ranking. A large positive PA Gap indicates that $f_{i}$ is truly spurious for $t$, as the model recognizes $t$ better when the spurious feature is present. In the object hallucination case, we apply the same methodology but to images that do not contain $t$. Here, a positive HR Gap means that the model hallucinates the presence of $t$ more frequently when $f_{i}$ is present, indicating that $f_{i}$ acts as a spurious cue that triggers hallucination. After computing the Gap for all potential spurious features, we select the feature with the largest Gap as the most influential spurious cue.
Validation To validate that SpurLens produces meaningful rankings, we quantatively compare SpurLens rankings against the original HardImageNet rankings, and against a baseline ranking method that takes the maximum of several random rankings. We find that SpurLens consistently outperforms both HardImageNet and the random baseline, demonstrating its effectiveness in detecting significant and interpretable spurious cues. Details of these validation experiments and results are found in Appendix I.
@@ 5 Results
Qwen2-VL 99.1 93.2 5.9 98.3 92.4 5.9 94.5 81.5 12.9
Llama-3.2 94.0 80.3 13.7 90.5 73.8 16.7 94.8 82.0 12.9
LLaVA-v1.6 95.6 83.2 12.4 93.2 78.3 14.9 95.3 81.5 13.7
GPT-4o-mini 95.9 83.2 12.6 94.9 83.6 11.3 87.6 69.7 17.9
Table 1: Object Recognition. Perception Accuracy PA (as %) across different datasets and models. $\text{PA}_{s}$ and $\text{PA}_{c}$ are accuracy on high-spurious and spurious-free images, respectively. All numbers are averaged over classes. When spurious cues are absent, all models struggle to recognize the target object.
Figure 4: The distribution of PA across ImageNet and COCO classes for Qwen2-VL. When spurious cues are absent, accuracy drops, and variance increases.
Models For our experiments, we evaluated three open-source MLLMs, Qwen2-VL-7B-Instruct [32], Llama-3.2-11B-Vision-Instruct [16], and LLaVA-v1.6-mistral-7B [10], along with one closed-source model, GPT-4o-mini [22]. All open-source models were accessed through HuggingFace.
Datasets We use three image datasets in our evaluation: HardImageNet [18], ImageNet [25], and COCO [9]. To ensure comparability with prior work, we use a subset of 100 classes from ImageNet, following the selection in Spurious ImageNet [20]. Each dataset offers diverse visual contexts, enabling us to analyze spurious correlations across different domains. The datasets also vary in complexity: HardImageNet is specifically curated to highlight spurious correlations, while COCO contains diverse real-world scenes, providing a broader evaluation of MLLM robustness.
Evaluation For each class in all three datasets, we initially generate 32 potential spurious features (see Appendix A). However, because we search within a single class and dataset sizes are limited, certain spurious features may not appear frequently enough. To ensure the reliability of our results, after applying the object detector, for each class we manually inspect the results for all spurious features to ensure that the top-$K$ and bottom-$K$ images do / do not contain the spurious cue and that these are correctly identified; those that fail this test are discarded. Note that this makes our reported Spuriosity Gaps more conservative. for each dataset, after computing the PA/HR Gaps, we take the class-wise average across the spurious features chosen by SpurLens and report the final aggregated metrics.
5.1 Object Recognition
The results are presented in Table 1. We use $K=50$ for HardImagenet and the ImageNet subset, and $K=100$ for COCO. We observe that performance decreases across all models when spurious cues are absent, with the magnitude of this drop varying between models and datasets. GPT-4o-mini exhibits the largest Gap on COCO, suggesting that it relies more on spurious correlations for object recognition compared to the other models. Some qualitative examples of GPT are provided in Appendix D. The larger performance drops in COCO can be attributed to its greater scene diversity and richer contextual information, which likely reinforce the model’s reliance on spurious cues. Furthermore, we provide qualitative examples in Appendix K, comparing model behavior with and without spurious cues. While the model generates detailed and accurate descriptions when spurious cues are present, it becomes less accurate when they are absent.
We conduct a class-wise analysis of spurious bias: the distribution of PA across ImageNet and COCO classes for Qwen2-VL [32] is visualized in Figure 4 (see Appendix J for other models). When spurious cues are absent, accuracy decreases and variance increases, indicating that models struggle to generalize without these cues. The distributions also confirm that spurious bias is highly class-dependent. We also performed a cross-model comparison. Figure 5 reveals that models exhibit inconsistent behavior across object classes. For instance, Qwen2-VL [32] is largely unaffected by spurious bias for snowmobile, whereas LLaVA-v1.6 [11] is highly impacted. This variability suggests that spurious correlations are not universally learned across models, in contrast to unimodal image classifiers, where biases tend to be more consistent [19]. We provide a detailed breakdown of the strongest spurious cues and their corresponding Spuriosity Gaps, as identified by SpurLens, for each model and class in Appendices L, M and N.
Figure 5: Comparing Normalized PA Gap Across Models. The heatmap shows that spurious bias affects different models inconsistently, with some objects being highly susceptible in certain models but not in others.
SpurLensTarget Object: BirdImage DatasetMLLM005555101010101515151520202020FeederNestPondSeedGrassy FieldHR Gap
@@ 5.1 Object Recognition
Yes, there is a bird in the image. It appears to be a small bird, possibly a type of parrot, perched on a feeder that is mounted on a tree…
@@
Yes, there is a bird in the image. It is a large bird, possibly a vulture or a similar species, with a prominent bald head and a dark body…
@@
Yes, there is a bird in the image. It is a small bird, possibly a parrot, located in the top left corner of the image. It is inside a cage that…
Figure 6: An example of hallucination identified by our method: (Left) SpurLens ranks the most influential spurious cues. (Right) The spurious cue, ‘feeder’ in this case, triggers hallucination in the model.
5.2 Seeing What’s Not There
We visualize an example of a spurious cue leading to hallucination, as identified by our method, in Fig. 6. Given an MLLM and a large image dataset, SpurLens detects the most influential spurious cues based on their Spuriosity Gap. For instance, for LLaVA [12] with target object “bird,” SpurLens identifies “feeder” as the most influential spurious cue. In Fig. 6, we demonstrate that this cue can indeed trigger hallucinations in the model. We provide more qualitative examples in Appendix C.
We use COCO and ImageNet as our datasets in this section; however, recall that to compute HR Gaps with SpurLens, for each class we require an image dataset that does not contain the target object. COCO is a large-scale object detection and segmentation dataset, whose 80 categories are grouped into 10 supercategories. For each target object, we run SpurLens on images from the same supercategory that do not contain the target object. We exclude three COCO classes (keyboard, dining table, and sports ball) from our final analysis because these objects frequently appear in the dataset without being annotated, making HR measurements unreliable for these categories. For ImageNet, for each of the 100 classes, we randomly sample 5,000 images outside of the class to serve as our dataset. We compute rankings with $K=50$ for ImageNet and $K=100$ for COCO. As an ablation, we provide an alternative setup using these two datasets in Appendix G.
The quantitative results are presented in Table 2. We observe that for both COCO and ImageNet, the hallucination rate is significantly higher when spurious cues are present. In COCO, the class-wise averaged HR is at least five times higher in high-spurious images across all models. In ImageNet, the effect is even more pronounced, with it being at least 13 times higher. These results demonstrate that spurious bias makes MLLMs vulnerable to hallucination, as the models fail to distinguish between true object presence and misleading contextual cues. Notably, GPT-4o-mini exhibits the lowest HR and the smallest gap, suggesting that it is less prone to hallucinating due to spurious bias compared to other models.
We conduct a class-wise analysis, similar to PA, to examine hallucination tendencies across different objects. The distribution of HR for Qwen2-VL is visualized in Fig. 7 (see Appendix J for other models). When spurious cues are present, hallucinations become significantly more pronounced. HR varies across classes, indicating that certain objects are more susceptible to hallucination than others. We provide detailed class-wise results, including the strongest identified spurious cue for each class, in Appendices L, M and N.
@@ 5.2 Seeing What’s Not There
Qwen2-VL 11.1 0.6 10.5 11.7 1.6 10.1
Llama-3.2 8.3 0.6 7.7 16.1 3.0 13.1
LLaVA-v1.6 18.8 1.4 17.4 20.4 2.8 17.6
GPT-4o-mini 4.6 0.3 4.3 9.8 1.1 8.8
Table 2: Hallucination Rates (HR as %) across different models and datasets. All numbers are averaged over classes. When spurious cues are present, the hallucination rate increases an order of magnitude.
Figure 7: The distribution of HR across ImageNet and COCO classes for Qwen2-VL. When spurious cues are present, hallucination amplifies.
6 Ablation Studies
6.1 Dropping Visual Cues
Both HardImageNet [18] and COCO [9] have pixel-level mask annotation of the target object. We leverage these to create artificial negative samples by removing visual cues of the target object and and evaluate the extent to which the model hallucinates its presence.
All open-source MLLMs have a vision encoder that patchifies the image, tokenizes each patch, and processes the resulting visual tokens. We drop the tokens associated with the target object just before they are processed by the vision encoder [3]. Since different MLLMs have varying architectures, we implement this token-dropping approach differently for each model; see Appendix F for implementation details. We also test hallucination rates on blank (fully black) images, where all visual cues are removed.
The results for this experiment are shown in Tab. 3. We observe significantly higher HR rates compared to Tab. 2. This poor performance likely arises because these images are out-of-distribution (OOD) samples, making them more challenging for the models. However, even under these OOD conditions, we observe that HR remains higher when spurious cues are present, reinforcing that models strongly rely on spurious correlations when the target object is absent. We also find that $\text{HR}_{b}$ is not always zero, meaning some models hallucinate objects even when given a completely blank input. These highlight that spurious bias in MLLMs is not solely dependent on visual features but is also reinforced by learned priors, making it a more fundamental issue than simple misclassification.
@@ 6.1 Dropping Visual Cues
Qwen2-VL 50.2 37.2 6.6 33.6 14.9 5.5
Llama-3.2 31.5 20.4 4.9 34.1 14.9 4.0
LLaVA-1.6 45.7 34.6 0 38.6 16.5 0
Table 3: Hallucination rates (HR as %) when visual cues of the target objects are removed through token-dropping. All numbers are averaged over classes.
@@ 6.2 Mitigation
Prompting can significantly influence how MLLMs behave. The Chain of Thought (CoT) prompting technique encourages the language model to generate rationales for their solution to a base prompt. Zero-shot CoT has been shown to improve performance on various benchmarks in both textual and multimodal scenarios [34, 6, 15, 36].
@@
In previous sections, we argue that these models tend to over-rely on spurious correlations. Now, we explore whether alternative prompting techniques can help mitigate this bias. We evaluate three strategies: (1) Prompt Ensembling, where we consider the majority vote among multiple queries with different prompt phrasing; (2) Dual Prompting, where we first ask the model for a general image description before prompting for object recognition; and (3) Attention-Guiding Prompting, where we explicitly instruct the model to focus on key details while avoiding misleading background cues. The prompts used in these strategies are provided in Appendix A.
The results are provided in Table 4. Perception Accuracy (PA) is evaluated on HardImageNet, while Hallucination Rate (HR) is evaluated on Spurious ImageNet. Additionally, we report $\text{HR}_{b}$, which measures hallucination rates on blank (fully black) images. While $\text{HR}_{b}$ is not always zero in the Baseline setting, it drops to zero with Ensemble or Dual Prompting, suggesting that these strategies help suppress some hallucinations.
Our results show a trade-off between accuracy and hallucination rates across different strategies. Dual Prompting reduces HR, but at the cost of lower PA, while Guiding Prompting improves PA but increases hallucinations, suggesting that encouraging the model to focus on key details might reinforce reliance on spurious cues. In contrast, Prompt Ensembling provides a minor but consistent improvement across all cases, reducing HR while slightly increasing PA, making it the most balanced strategy. However, despite these variations, the Spuriousity Gap remains present across all prompting strategies, suggesting that while prompt engineering can influence model behavior, it does not eliminate the underlying reliance on spurious cues. This indicates that the spuriousity gap is a more fundamental issue that cannot be resolved solely through prompting techniques.
Model Prompt $\text{PA}_{s}$ $\text{PA}_{c}$ $\text{HR}_{s}$ $\text{HR}_{c}$ $\text{HR}_{b}$
Qwen2 Baseline 98.1 91.9 9.5 1.2 3.3
Ensemble 98.5 92.4 8.6 1 0
Guiding 97.5 89.3 16.4 0.8 0
Dual 95.2 82.3 4 0.4 0
Llama Baseline 95.1 84.7 11 2.2 7
Ensemble 96.8 87 7.5 1.5 0
Guiding 95.6 87 15.1 4.1 4.7
Dual 95.2 84.4 7.6 3.1 0
LLaVA Baseline 91.4 85.2 12.5 2.1 0
Ensemble 92.4 86.4 11 1.9 0
Guiding 99.2 91.2 24.6 5.8 0
Dual 89.3 72 5.6 0.8 0
@@
Table 4: Mitigating Spurious Bias. Perception Accuracy (PA%) and Hallucination Rate (HR%) across different prompting strategies and models. PA is measured on HardImageNet, while HR is measured on Spurious ImageNet. Ensemble improves the Baseline in all cases.

### v2

@@ 4 SpurLens: Automating Discovery of Spurious Cues
$\textbf{PA}_{s}=\frac{1}{K}\sum\nolimits_{\mathcal{I}\in\mathcal{U}_{t,f_{i}}^{+}}\textbf{PA}(\mathcal{M},\mathcal{I},t),\;\textbf{PA}_{c}=\frac{1}{K}\sum\nolimits_{\mathcal{I}\in\mathcal{U}_{t,f_{i}}^{-}}\textbf{PA}(\mathcal{M},\mathcal{I},t),\;\textbf{PA Gap}=\textbf{PA}_{s}-\textbf{PA}_{c}$ (6)
In the object recognition case, the Spurious Gap measures the difference in recognition accuracy between images with and without $f_{i}$, based on the top-$K$ and bottom-$K$ ranked images in the $f_{i}$-score ranking. A large positive PA Gap indicates that $f_{i}$ is truly spurious for $t$, as the model recognizes $t$ better when the spurious feature is present. In the object hallucination case, we apply the same methodology but to images that do not contain $t$. Here, a positive HR Gap means that the model hallucinates the presence of $t$ more frequently when $f_{i}$ is present, indicating that $f_{i}$ acts as a spurious cue that triggers hallucination. After computing the Gap for all potential spurious features, we select the feature with the largest Gap as the most influential spurious cue.
Validation In Appendix E, we compare GPT-4’s spurious feature proposals with other LLMs (e.g. Gemini) and observe substantial overlap, suggesting that our results are not dependent on the choice of language model. In Appendix I, we study the sensitivity of the computed Gaps to the choice of $K$, and find that the Gaps stabilize for all models as $K$ increases, indicating the SpurLens results are robust to this parameter. In Appendix F, we assess the accuracy of the OWLv2 object detector with two human studies. First, we randomly sample detection outputs and check whether human annotators agree with the presence/absence of the spurious cue in the top- and bottom-ranked images; we find high agreement: 93% on COCO and 90% on ImageNet. Second, we for a random subset of spurious cues identified by SpurLens, we replace the object detector results with human annotations, and compute Gaps based on the human-labeled top- and bottom-ranked images; we observe strong alignment with the SpurLens gaps. This demonstrates that, while the object detector may itself be subject to some spurious biases, our methodology is sufficiently robust to enable reliable detection of large spurious gaps. In Appendix N, we compare SpurLens gaps to a random baseline, providing evidence that the signals SpurLens detects are statistically significant.
Finally, we note that SpurLens emphasizes precision, rather than recall or completeness: it aims to consistently and robustly find strong spurious cues, rather than the absolute best cue for each model and dataset. While our results cannot be used to directly compare models, we do find significant spurious biases in all models that we evaluate. Although SpurLens uses binary-resemblance prompts to evaluate MLLMs (for ease of evaluation and minimizing other sources of bias), we find many instances of the identified biases manifesting in open-ended generation (see Appendices C, D, and R).
@@ 5 Results
Models For our experiments, we evaluate three open-source MLLMs, Qwen2-VL-7B-Instruct (Qwen2VL), Llama-3.2-11B-Vision-Instruct (llama3), and LLaVA-v1.6-mistral-7B (improvedllava), along with one closed-source model, GPT-4o-mini (openai2024gpt4omini).
Datasets We use three image datasets in our evaluation: HardImageNet (hardimagenet), ImageNet (imagenet), and COCO (coco). To maintain comparability with prior work, we use a subset of 100 classes from ImageNet as chosen in Spurious ImageNet (spnet). These datasets offer diverse visual contexts, enabling analysis of spurious correlations across different domains. They also vary in complexity: HardImageNet is curated to highlight spurious correlations, while COCO contains diverse real-world scenes, providing a broader evaluation of MLLM robustness. In contrast, older benchmarks such as CelebA (celeba) and Waterbirds (waterbirds) were designed for unimodal classifiers trained end-to-end, and thus fail to challenge MLLMs in a zero-shot setting. See Appendices G and H for further discussion.
Evaluation For each class in all three datasets, we initially generate 32 potential spurious features, which are then passed through various filters (see Appendix A). We compute the PA/HR Gap for each feature, and choose the most influential one as the SpurLens feature for that class. For each dataset, we take the class-wise average of the chosen PA/HR Gaps and report the final aggregated metrics.
5.1 Object Recognition
The results are presented in Table 1. We use $K=50$ for HardImageNet and the ImageNet subset, and $K=100$ for COCO. We observe that performance decreases across all models when spurious cues are absent, with the magnitude of this drop varying between models and datasets. Some qualitative examples of GPT-4o-mini failures are provided in Appendix D. Additional qualitative examples are found in Appendix R, which compares the behavior of open-source models with and without spurious cues. While the models generate detailed and accurate descriptions when spurious cues are present, the descriptions become less accurate when they are absent.
We conduct a class-wise analysis of spurious bias: the distribution of PA across ImageNet and COCO classes for Qwen2-VL (Qwen2VL) is visualized in Figure 4 (see Appendix Q for other models). When spurious cues are absent, accuracy decreases and variance increases, indicating that models struggle to generalize without these cues. The distributions also confirm that spurious bias is highly class-dependent. We provide a detailed breakdown of the strongest spurious cues and their corresponding Spurious Gaps as identified by SpurLens, for each model and class, in Appendix S.
5.2 Object Hallucination
We visualize an example of a spurious cue leading to hallucination, as identified by SpurLens, in Figure 3. Given an MLLM and a large image dataset, SpurLens detects the most influential spurious cues based on their Spurious Gap. For LLaVA (llavanext) with target object “bird”, SpurLens identifies “feeder” as the most influential spurious cue. In Figure 3, we demonstrate that this cue can indeed trigger hallucinations in the model. We provide more qualitative examples in Appendix C.
Recall that to compute HR Gaps with SpurLens, for each class we require an image dataset that does not contain the target object. COCO is a large-scale object segmentation dataset, whose 80 categories are grouped into 10 supercategories. For each target object, we run SpurLens on images from the same supercategory that do not contain the target object. We exclude three COCO classes (keyboard, dining table, and sports ball) from our final analysis because these objects frequently appear in the dataset without being annotated, making HR measurements unreliable for these categories. For ImageNet, for each of the 100 classes, we randomly sample 5,000 images outside of the class to serve as our dataset. We compute rankings with $K=50$ for ImageNet and $K=100$ for COCO. As an ablation, we provide an alternative setup using these two datasets in Appendix L.
The quantitative results are presented in Table 2. We observe that for both COCO and ImageNet, the hallucination rate is significantly higher when spurious cues are present. In COCO, the class-wise averaged HR is at least 5 times higher in high-spurious images across all models; the effect is more pronounded in ImageNet, with it being at least 18 times higher. These results demonstrate that spurious bias makes MLLMs vulnerable to hallucination, as models fail to distinguish between true object presence and misleading contextual cues.
Similar to PA, we conduct a class-wise analysis, to examine hallucination across objects. The distribution of HR for Qwen2-VL is visualized in Figure 4 (see Appendix Q for other models). When spurious cues are present, hallucinations become significantly more pronounced. HR varies across classes, indicating that some objects are more susceptible to hallucination than others. We provide detailed class-wise results, in Appendix S.
Table 1: Object Recognition. Perception Accuracy (PA%) across different datasets and models. $\text{PA}_{s}$ and $\text{PA}_{c}$ are accuracy on high-spurious and spurious-free images. Results are averaged over classes.
@@ 5.2 Object Hallucination
Qwen2-VL 99.3 93.4 5.9 98.4 91.7 6.7 95.3 80.1 15.2
Llama-3.2 93.1 78.1 15.0 90.6 72.9 17.7 95.0 80.0 14.8
LLaVA-v1.6 96.0 82.8 13.2 93.0 77.3 15.7 95.6 79.6 16.0
GPT-4o-mini 95.0 82.4 12.6 95.7 83.8 11.9 88.0 67.6 20.4
SpurLensTarget Object: BirdMLLM055101015152020FeederNestPondSeedFieldHR Gap
@@
Yes, there is a bird in the image. It appears to be a small bird, … perched on a feeder that is mounted on a tree…
@@
Yes, there is a bird in the image. It is a large bird, possibly a vulture … with a prominent bald head and a dark body…
@@
Yes, there is a bird in the image. It is a small bird, … in the top left corner of the image. It is inside a cage that…
Figure 3: An example hallucination identified by SpurLens: (Left) SpurLens ranks the most influential spurious cues. (Right) The spurious cue, ‘feeder’ in this case, triggers hallucination in the model.
6 Ablation Studies
6.1 Dropping Visual Cues
Both HardImageNet (hardimagenet) and COCO (coco) have pixel-level mask annotation of the target object. We leverage these to create artificial negative samples by removing visual cues of the target object and and evaluate to what extent the models hallucinate their presence.
All open-source MLLMs have a vision encoder that patchifies the image, tokenizes each patch, and processes the resulting visual tokens. We drop the tokens associated with the target object before they are processed by the vision encoder (tokendropping). Since different MLLMs have varying architectures, we implement this token-dropping approach separately for each model; see Appendix K for implementation details. We also test hallucination rates on blank (fully black) images, where all visual cues are removed.
The results for this experiment are shown in Table 3. We observe significantly higher HR rates compared to Table 2. This poor performance likely arises because these images are out-of-distribution (OOD) samples, making them more challenging for the models. However, even under these OOD conditions, we observe that HR remains higher when spurious cues are present, reinforcing that models strongly rely on spurious correlations when the target object is absent. We also find that $\text{HR}_{b}$, the hallucination rate on blank (fully black) images, is not always zero, meaning some models hallucinate objects even when given a completely blank input. This highlights that spurious bias in MLLMs is not solely dependent on visual features.
Figure 4: Per-class PA and HR distributions for Qwen2-VL across two datasets.
Table 2: Hallucination rates (HR %) for high-spurious vs spurious-free images. When spurious cues are present, hallucination rates increase.
@@ 6.1 Dropping Visual Cues
Qwen2-VL 12.1 0.5 11.6 13.5 1.8 11.7
Llama-3.2 9.0 0.5 8.6 18.1 3.8 14.4
LLaVA-v1.6 18.4 1.0 17.4 21.8 2.7 19.2
GPT-4o-mini 5.2 0.2 5.1 11.2 1.4 9.8
Table 3: Hallucination rates (HR %) after object token-dropping. Models hallucinate even when target object features are removed.
@@
Qwen2-VL 50.2 37.2 6.6 35.3 15.1 5.5
Llama-3.2 31.5 20.4 4.9 34.8 15.1 4.0
LLaVA-1.6 45.7 34.6 0.0 40.0 16.4 0.0
@@ 6.2 Mitigation
SpurLens is developed as a diagnostic tool that automatically identifies interpretable causes of spurious bias, which can be used for mitigation. However, in this section, we establish that the spurious biases identified are a fundamental issue and cannot be trivially overcome through language-based techniques. We focus on Chain-of-Thought (CoT) prompting, which encourages language models to generate rationales for their solution to a base prompt. Zero-shot CoT has previously been shown to improve performance on various benchmarks in both textual and multimodal scenarios (wei2023chainofthoughtpromptingelicitsreasoning; cot; luo2024pkrdcotunifiedchainofthoughtprompting; zhang2024multimodalchainofthoughtreasoninglanguage).
@@
We evaluate five strategies. (1) Prompt Ensembling, where we take the majority vote among multiple queries with different prompt phrasing. (2) Dual Prompting, where we first ask for a general image description before prompting for object recognition. (3) Guiding Prompting, where we explicitly instruct the model to focus on key details and avoid misleading background cues. (4) Spurious List, where we provide the model with a list of potential spurious features (the list of GPT-4 suggested features that SpurLens analyzes). (5) Spurious Top, where we provide the model with the strongest spurious feature as identified by SpurLens (note that this is improper, as it incorporates SpurLens’ knowledge of the full dataset). The prompts used in these strategies are provided in Appendix A.
The results are provided in Table 4. PA is evaluated on HardImageNet, while HR is evaluated on Spurious ImageNet (spnet). While the ensemble/reasoning-based strategies offer slight improvements on certain models and metrics, none of the strategies perform significantly better or have substantially smaller Spurious Gap compared to the baseline. The Spurious Top strategy “cheats” compared to the other strategies by using SpurLens’ analysis and identification of the strongest spurious cue on each class’ dataset, which is the same dataset the model is being evaluated on; yet, it also does not perform well relatively. Further discussion of these strategies and their relative performance can be found in Appendix J. The continued presence of the Spurious Gap across all strategies suggests that language-based methods, while influencing the model’s behavior, cannot eliminate the underlying reliance on spurious cues; Spurious Gaps detected by SpurLens are a more fundamental issue that cannot be resolved solely through prompting techniques.
@@
Table 4: Mitigation Attempt via Prompting. Perception Accuracy (PA%) and Hallucination Rate (HR%) across different prompting strategies and models. The Spurious Gap persists in all strategies.
Model Metric Baseline Ensemble Guiding Dual Spur. List Spur. Top
Qwen2 $\text{PA}_{s}$ 98.1 98.5 97.5 95.2 96.5 97.0
$\text{PA}_{c}$ 91.9 92.4 89.3 82.3 85.9 88.9
$\text{HR}_{s}$ 9.5 8.6 16.4 4.0 5.5 4.8
$\text{HR}_{c}$ 1.2 1.0 0.8 0.4 0.5 0.8
$\text{HR}_{b}$ 3.3 0 0 0 0 1.0
Llama $\text{PA}_{s}$ 95.1 96.8 95.6 95.2 73.3 95.2
$\text{PA}_{c}$ 84.7 87.0 87.0 84.4 63.2 85.6
$\text{HR}_{s}$ 11.0 7.5 15.1 7.6 4.8 10.3
$\text{HR}_{c}$ 2.2 1.5 4.1 3.1 1.0 2.0
$\text{HR}_{b}$ 7.0 0 4.7 0 0.7 0.6
LLaVA $\text{PA}_{s}$ 91.4 92.4 99.2 89.3 94.8 89.6
$\text{PA}_{c}$ 85.2 86.4 91.2 72.0 81.2 82.7
$\text{HR}_{s}$ 12.5 11.0 24.6 5.6 11.1 9.5
$\text{HR}_{c}$ 2.1 1.9 5.8 0.8 1.6 1.4
$\text{HR}_{b}$ 0 0 0 0 0 0